OpenStudy (anonymous):

Solve using the addition method. What is the y-value of the solution to the system? 2x - 3y = -1 3x + 4y = 24

OpenStudy (texaschic101):

the addition method, I believe, is the same as the elimination method 2x - 3y = -1 ---->(3) 3x + 4y = 24 --->(-2) ------------------- 6x - 9y = -3 (result of multiplying by 3) -6x - 8y = -48 (result of multiplying by -2) -------------------add -17y = - 51 y = -51/-17 y = 3 2x - 3y = -1 2x - 3(3) = -1 2x - 9 = -1 2x = 9 - 1 2x = 8 x = 4 check.. 3x + 4y = 24 3(4) + 4(3) = 24 12 + 12 = 24 24 = 24 (correct) I know you just needed the y term, but I went ahead and solved for x as well, so I could check my answers. Answer is : y = 3